OpenStudy (anonymous):

help please. This is confusing to me. sqrt of x-m = sqrt of x-5

OpenStudy (doc.brown):

\[\sqrt{x-m}=\sqrt{x-5}\]\[\sqrt{x-m}^2=\sqrt{x-5}^2\]\[x-m=x-5\]\[-m=-5\]\[m=5\]

OpenStudy (doc.brown):

It kind of makes more sense if you just look at it.\[\sqrt{x-m}=\sqrt{x-5}\]Clearly m=5 if both sides of the equal sign are equal.

OpenStudy (anonymous):

wow.. right. But i was trying to get in track of how you work it out but yes i shouldve just see it. but thank you very much

OpenStudy (doc.brown):

You're welcome. I squared both sides to remove the square root. Subtracted x from both sides. Then multiplied both sides by -1. That kept both sides equal all the times... which they have to be.

OpenStudy (anonymous):

kk.. thank you. I love you.. as a helping little owl

OpenStudy (doc.brown):

I love you too.